OpenStudy (anonymous):

Form a polynomial function whose real zeros and degree are given. Zeros: -3, 0, 4; degree 3

OpenStudy (john_es):

A polinomy with n degree and zeroes at x0, x1, x2,...,xn is \[(x-x_0)(x-x_1)\cdots(x-x_n)\]So, \[(x+3)(x-0)(x-4)=x(x+3)(x-4)=x^3-x^2-12 x\]